OpenStudy (anonymous):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle? (x - 1)2 + (y + 2)^2 = 3 (x + 1)2 + (y - 2)^2 = 3 (x + 1)2 + (y + 2)^2 = 9 (x + 1)2 + (y - 2)^2 = 9

OpenStudy (anonymous):

(x-(-1))^2+(y-2)^2=3^2 (x+1)^2+(y-2)^2=9

OpenStudy (anonymous):

d is answer